OpenStudy (anonymous):

ordered pairs (1,2),(2,4),(3,8) what is the function equation

OpenStudy (anonymous):

if i was going to guess i would say \(f(x)=2^x\) since \(2^1=2, 2^2=4, 2^3=8\)

OpenStudy (anonymous):

Wow that was easier than I thought. It makes perfect sense now. Thanks so much!